OpenStudy (anonymous):

Ferdinand wrote a check for $96 to pay his monthly cable bill, but when balancing his checkbook, he accidentally recorded it as a credit rather than as a debit?

jimthompson5910 (jim_thompson5910):

What's your question?

OpenStudy (anonymous):

how will his check register compare to his monthly bank statment when he receives it?

jimthompson5910 (jim_thompson5910):

Well he'll have 2*96 = 192 dollars more than he actually does

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

So say his balance is $1000 He writes the check for $96. He *should* have adjusted his balance to 1000 - 96 = 904 dollars However, he added 96 to his balance by mistake making it 1000 + 96 = 1096 dollars

jimthompson5910 (jim_thompson5910):

So in his check register, he'll have 1096 as the balance but the bank statement will be 904 This is a 1096 - 904 = 192 dollar difference

OpenStudy (anonymous):

Thank you! that helped!!

jimthompson5910 (jim_thompson5910):

you're welcome